# Random Stability of Random Variables

### 30 Mar 2026 v1

@@ 1.1. The main definition
Alternatively, the article [40] uses the term branching stability to stress connection with branching processes. We will see in our article that this connection is indeed critical. In [24, Definition 3], the term strictly $N$-stable is used.
@@ 1.2. Motivation
For random stability, there has been a lot of research starting from the 1990s, as well as some earlier research in relation to the classic theory of branching processes. But existing literature is fragmented. Equivalent formulations have appeared in difference forms, and this article aims to unify and extend them. It contains exposition, new research, and critique of existing research. Theorems 1,  2, and 3 are our main results. We could not find Theorem 1 in other articles. Theorem 2 was shown in [32] but under special assumptions. Theorem 3 was shown in special cases, see [13].
@@ 2.1. Existence results
Theorem 1 is the main result of this article. As mentioned earlier, we could not find it in the literature, despite its importance. The closest was [37, Corollary 1.2] with (in their notation) $A_{i}=1/c$. This discusses a generalization of random stability to so-called it smoothing problems. But their proof applies only to nonnegative $N$-stable $X$.
@@
There are two cases, when $c=\mathbb{E}[N]\in(1,\infty)$; see also discussion in [5]. For a conceptual proof, see [39]. For other proofs, see [49, Theorem 3.1], [50, Theorem 4.3].
@@ 2.2. Strictly stable distributions on the real line
Recall a well-known classification of stictly stable distributions $Q$, see [47, Chapter 1] or [53, Chapter V]. Its characteristic function (Fourier transform)
@@
A standard reference is the classic monograph [47, Chapter 1]: Equivalent Definitions 1.1.1, 1.1.4, 1.1.5, 1.1.6 of stable distributions, and Properties 1.2.6, 1.2.8 of strictly stable distributions. Also, see other monographs: [38, Chapter 3]: Theorem 5.7.3 and discussions after this; a more recent book [53, Chapter V]: Section 7 for general results about strict stability; Theorem 7.6 for strictly stable symmetric distributions; Theorem 3.5 for positive strictly stable symmetric distributions; Theorem 3.5 for positive strictly stable random variables. Finally, we point out to [13, Chapter 4], section 4.5, page 130. We stress that much of classic research involves stable not strictly stable distributions, which arise as weak limits
@@ 2.3. Classification of $N$-stable $X$ for given NN
In Theorem 2, we classify $N$-stable distributions. This replicates the results [32, Proposition 2.1, Proposition 3.1] but without an artificial additional assumption in Remark 2 below. The $N$-stable laws are represented via a product involving a standard $N$-stable random variable, and an independent strictly $\alpha$-stable random variable.
@@ 2.4. Removal of artificial assumptions
This condition was first introduced in [55] and used for the converse of the transfer theorem. This theorem is about a transfer from weak limits of sums of independent random variables to weak limits of random sums of these variables, [13, Chapter 4]: Section 4.1, Theorem 4.1.2, and references therein. Says [55]: The following condition is presumably true, but so far no one was able to prove or disprove it. It turns out from our research that we do not, in fact, need to prove this condition! Indeed, from 1973 until now, to the best of our knowledge, no one could prove this in the general form. There were many discussions of this condition in a recent book [13], which is an unabridged republication from 1996: [13, Chapter 4], Section 2. To quote: No one has managed either to prove this or to refute it yet. There, it was proved in several cases. In [22, Chapter 8], it was mentioned but unproved in Remark 8.8.16.
@@ 2.5. Examples of random stable variables
We stress its similarity with classic Gaussian random variables. In the literature, sometimes a Gaussian $N$-stable $X$ is defined using (1) with (9) with $\alpha=2$. An additional restriction $\mathbb{E}[X^{2}]=\infty$ is introduced in [23, Definition 2]. Confusingly, the opposite restriction $\mathbb{E}[X^{2}]<\infty$ was introduced in [13, Chapter 4], see Definition 4.6.1; or [24, Definition 2], or [22, Chapter 8], Definition 8.3.2. But we do not need these restrictions in the definition of a Gaussian $N$-stable random variable, as seen in the following lemma.
@@ 2.6. Geometric stability
with $0<\alpha\leq 2$. See also [22, Chapter 9], Subsection 9.4.2.1, Theorem 9.4.6. This distribution is absolutely continuous on the real line, but its density is explicitly known only for $\alpha=2$, when it becomes the Laplace distribution. The positive $N_{p}$-stable distribution has Laplace transform with $\beta>0$ and $\alpha\in(0,1]$:
@@
This is commonly known as the Mittag-Leffler distribution; see [22, Chapter 9], Subsection 9.4.2.2, Theorem 9.4.9. Also, in [13, Chapter 2], these were named Kovalenko distributions, see (3.2) or page 30. This distribution is also absolutely continuous, but its density is explicitly known only for $\alpha=1$ or $\alpha=0.5$. In the former case, this distribution is exponential, as in (13). In the latter case, for $\beta=1$, it has density
@@
These were the earliest and most studied examples. An entire chapter [22, Chapter 9] is devoted to these geometric stable distributions. See also a great survey [29] with large bibliography. Such geometric stable distributions were studied as far back as in 1984 in [25] and also in one early article by T. Kozubowski [28]. Further representation results for the Linnik and the Mittag-Leffler distributions are given in [27].
@@ 2.7. The infinitely divisible case
The following example is well known, see [17, Chapter I, Example 8.5; Chapter V, Example 13.2], [40, Section 8], [23, Section 6], Exampples 2 and 4, [7, Example 1]; [44, Example 3], and physical applications in [16, Section 8]. Here $X$ is a Gamma random variable with shape $1/k$ and certain rate parameter $\beta>0$. We remind the readers that it has Laplace transform
@@ 2.8. Commuting distributions
The key result is below. It was shown in other literature in particular cases, but not in this general form. For such results for or Gaussian $N$-stable random variables, see [13, Chapter 4], Theorem 4.6.1 and Corollary 4.6.1; [24, Theorem 1], [22, Chapter 8], Theorem 8.3.4 and references therein; [44, Theorem 2.1]. The property of commutative semigroups was also shown in [29, Remark 2.4]. This survey [29] is very comprehensive for 1990s. However, they claim that geometric family is unique among such explicit commuting families. We discuss in Section 2 that this is not true.
@@ 3.1. Introduction to the problem
The idea of the proof of Lemma 7 is to consider the behavior of the probability generating function of $N$ at $s=0$. From (2), which involves the behavior of the Laplace transform of $X$ at infinity. This approach is useful to disprove that $\varphi$ is a probability generating function in other cases. Such distribution $Q$ might have an atom at 0, that is, have $Q({0})=\mathbb{P}(X=0)>0$; and this happens if and only if $N$ has an atom at zero: $\mathbb{P}(N=0)>0$. By the way, the case of an atom at zero can be reduced to the case of no atom at zero. See [1, Chapter I], Section 12; [40, Section 9]; [16, Section 3], Definitions 3.1, 3.2; Theorem 3.2.
@@ 3.4. Parameterization
In the second case, we cannot have both expectation and convergence assumptions for $\nu_{p}$. The notation of $\nu_{p}$ is used in the following literature (the list is not exhaustive): [44, Section 2], (1) and (3), both expectation and convergence; [32, Section 1], only convergence; [23, page 304], they mentioned only expectation but clearly meant convergence as well, as seen from Theorem 2; [24, Section 2], only expectation, but later convergence is introduced; [32, Section 2] (3) and (4), only convergence; [13, Chapter 4], (6.32), only convergence. Sometimes, commutativity is required or derived in special cases.
The real reason behind the prevalence of this $p$-notation is that, historically, random stability study to a large extend (although not exclusively) was focused on geometric stability (that is, when $N=N_{p}$ is geometric from (11)). In this case, $pN_{p}$ converges to the standard exponential distribution as $p\downarrow 0$, and $\mathbb{E}[N_{p}]=1/p$. Approximations of sums of $N_{p}$ many independent identically distributed random variables as $p\downarrow 0$ (similar in spirit to the Central Limit Theorem) were extensively studied. A good exposition is in [22, Chapter 9].
@@ 4. Equivalent Definitions of Random Stability
There is another definition of random stable variables in the literature, one that is based on weak limits of scaled random sums. Just as stable and strictly stable distributions (in the classic sense) are often defined as weka limits of scaled sums of $N$ independent identically distributed random variables as $N\to\infty$, so can be done for random stable distributions. See [32, Section 2], and [40, Theorems 1, 2, 3].
@@
We see that $U_{j}$ belong to the strict domain of attraction: Recall our discussion above about defining strictly stable distributions as weak limits of scaled sums of independent identically distributed random variables. We stress this is different from classic $\alpha$-stable domains of attraction, where adding constants is allowed: Recall (6) and the discussion there. Strict domain of stability include scaling sums of independent indentically distributed random variables, but not adding constants. See also discussion in [32, Section 2].
@@ 5.3. Cyclic semigroup
$S(X)=\{1,a,a^{2},\ldots\}$ for some $a>1$. A construction was given in [7, Examples 2, 4] based on [9, Section 5] and [4] but in our view the proofs are not witten well, and they should be improved. Particular cases of $Q$ discussed in [4, Section 6] are as follows: The standard $P$-stable $Q$ for the following $P$: (a) $P(\{1\})=P(\{2\})=0.5$, with $\varphi(s)=(s+s^{2})/2$; (b) $P$ is the shifted geometric, with probability generating function $\varphi(s)=s^{2}/(4-3s)$, used in [2, Section 3] as a tool to construct Brownian motion on the Sierpinski gasket. The idea is to create an almost constant solution $\theta:(0,\infty)\to(0,\infty)$ to the functional equation $\theta(ax)\equiv b\theta(x)$, for fixed $a,b>0$. See [57] for additional information, also [48, Section 4].
@@ 5.5. Complete squares
$S(X)=\{1,4,9,16,\ldots\}$. This was discussed in [23, Section 4] for $L(u)=1/\mathrm{ch}(\sqrt{2u})$, where $\mathrm{ch}(x)=(e^{x}+e^{-x})/2$ is the hyperbolic cosine, with the inverse function $\mathrm{arcch}(s)=\ln(s+\sqrt{s^{2}-1})$.
@@
This equality in law follows from the self-similarity of the Brownian motion, see [20, Chapter 2]. We also refer the reader to [23, Section 3] with an interesting discussion about rational probability generating functions.
@@ 6.1. Construction
This branching process is described as follows, as outlined in [17, Chapter 5], Theorem 9.1, and [1, Chapter III], Sections 2 and 3, see also a short discussion in [53, Chapter V], Section 8. Start, as in discrete time, from $N(0)=1$. Fix a distribution $H$ on $\{0,2,3,\ldots\}$ called the generating distribution. This $H$ plays the role similar to the infinitesimal generator for diffusion processes. Each particle which exists at time $t$ is replaced during time interval $t,t+\triangle t]$ with small $\triangle t$, with probability $\triangle t$ with the number of particles distributed as $H$.
@@ 6.2. Yule and Neveu processes
For geometric distributions, $(N(t),\,t\geq 0)$ is called the Yule process, introduced by Yule in [56, Section 1] in 1925 for modeling evolution of the number of species as the pure birth process with birth rate $n$ at state $n$. See the discussion in [17, Chapter 5, Section 8] or [1, Chapter III], Section 5. For Sibuya distributions, this is called the Neveu process, introduced in an unpublished manuscript [46] by Neveu in relation to Generalized Random Energy Model (GREM) of Derrida. Links to Sherrington-Kirkpatrick model and GREM are discussed in [3, Chapter 6]. The exponential $\mathrm{Exp}(1)$ from (13) is $G_{p}$-stable for any $p$. However, there is no Sibyua-stable random variable, since the mean of the Sibyua distribution is zero. Properties of the Neveu process are discussed in [18]. They also considered generalizations with Sibuya distribution with an atom at zero in [18]. These branching processes have nonzero extinction probability.
@@ 6.3. Generating geometric distribution
Here we present a special case of geometric generating $H$. The full treatment is given in a recent article [42], with complicated formulas for mass functions and PGFs, using special functions. Our example is easily treatable and explicit. We also refer to another article [43] by the same authors, treating the case of Poisson $H$. This involves complicated special functions, too. Similar computations were done for [41, Theorem 1] by the same authors; but they consider the subcritical case $c<1$, with a.s. extinction; of course, this branching process does not have a limiting $X$. One possible application is to physics: the Bose-Einstein distribution have geometric as one possible limit, see [19, (9)], and this can be connected to the current process.
@@ 6.4. Theta-branching processes
In a recent article [36], they introduced a classification of 9 cases, see Section 4, with generalization of a Neveu process with Sibuya distribution. It has four parameters, including $\theta\in[-1,1]$. The case 4 is precisely the Sibuya distribution modified so that there is an atom at zero, and a nonzero extinction probability. For us, the case 3 is the most interesting, since it corresponds to the supercritical case, but with finite mean. The probability generating function $h$ of the generating distribution $H$ is, see [36, Section 7]:
@@
The probability generating function of the continuous-time branching process at time $t$ with $p=e^{-t}$, see [36, Section 4]:

### 02 Apr 2026 v2

@@ 1.1. The main definition
Alternatively, the article [38] uses the term branching stability to stress connection with branching processes. We will see in our article that this connection is indeed critical. In [25, Definition 3], the term strictly $N$-stable is used.
@@ 1.2. Motivation
For random stability, there has been a lot of research starting from the 1990s, as well as some earlier research in relation to the classic theory of branching processes. See [29] for a survey of geometric stable distributions ($N$-stable for geometric $N$), and a more detailed description in [23, Chapter 9]. A classic application for classic stable and random stable laws is asset returns, see [29, Remark 2.2] and references therein; also [22, Chapter 4], page 157. For $N$-stable random variables with general $N$, see [22, Chapter 4].
However, existing literature is fragmented. Equivalent formulations have appeared in difference forms, and this article aims to unify and extend them. It contains exposition, new research, and critique of existing research. Theorems 1,  2, and 3 are our main results. We could not find Theorem 1 in other articles. Theorem 2 was shown in [30] but under special assumptions. Theorem 3 was shown in special cases, see [13].
@@ 2.1. Existence results
Theorem 1 is the main result of this article. As mentioned earlier, we could not find it in the literature, despite its importance. The closest was [35, Corollary 1.2] with (in their notation) $A_{i}=1/c$. This discusses a generalization of random stability to so-called it smoothing problems. But their proof applies only to nonnegative $N$-stable $X$.
@@
There are two cases, when $c=\mathbb{E}[N]\in(1,\infty)$; see also discussion in [5]. For a conceptual proof, see [37]. For other proofs, see [47, Theorem 3.1], [48, Theorem 4.3].
@@ 2.2. Strictly stable distributions on the real line
Recall a well-known classification of stictly stable distributions $Q$, see [45, Chapter 1] or [51, Chapter V]. Its characteristic function (Fourier transform)
@@
A standard reference is the classic monograph [45, Chapter 1]: Equivalent Definitions 1.1.1, 1.1.4, 1.1.5, 1.1.6 of stable distributions, and Properties 1.2.6, 1.2.8 of strictly stable distributions. Also, see other monographs: [36, Chapter 3]: Theorem 5.7.3 and discussions after this; a more recent book [51, Chapter V]: Section 7 for general results about strict stability; Theorem 7.6 for strictly stable symmetric distributions; Theorem 3.5 for positive strictly stable symmetric distributions; Theorem 3.5 for positive strictly stable random variables. Finally, we point out to [13, Chapter 4], section 4.5, page 130. We stress that much of classic research involves stable not strictly stable distributions, which arise as weak limits
@@ 2.3. Classification of $N$-stable $X$ for given NN
In Theorem 2, we classify $N$-stable distributions. This replicates the results [30, Proposition 2.1, Proposition 3.1] but without an artificial additional assumption in Remark 2 below. The $N$-stable laws are represented via a product involving a standard $N$-stable random variable, and an independent strictly $\alpha$-stable random variable.
@@ 2.4. Removal of artificial assumptions
This condition was first introduced in [53] and used for the converse of the transfer theorem. This theorem is about a transfer from weak limits of sums of independent random variables to weak limits of random sums of these variables, [13, Chapter 4]: Section 4.1, Theorem 4.1.2, and references therein. Says [53]: The following condition is presumably true, but so far no one was able to prove or disprove it. It turns out from our research that we do not, in fact, need to prove this condition! Indeed, from 1973 until now, to the best of our knowledge, no one could prove this in the general form. There were many discussions of this condition in a recent book [13], which is an unabridged republication from 1996: [13, Chapter 4], Section 2. To quote: No one has managed either to prove this or to refute it yet. There, it was proved in several cases. In [23, Chapter 8], it was mentioned but unproved in Remark 8.8.16.
@@ 2.5. Examples of random stable variables
We stress its similarity with classic Gaussian random variables. In the literature, sometimes a Gaussian $N$-stable $X$ is defined using (1) with (9) with $\alpha=2$. An additional restriction $\mathbb{E}[X^{2}]=\infty$ is introduced in [24, Definition 2]. Confusingly, the opposite restriction $\mathbb{E}[X^{2}]<\infty$ was introduced in [13, Chapter 4], see Definition 4.6.1; or [25, Definition 2], or [23, Chapter 8], Definition 8.3.2. But we do not need these restrictions in the definition of a Gaussian $N$-stable random variable, as seen in the following lemma.
@@ 2.6. Geometric stability
with $0<\alpha\leq 2$. See also [23, Chapter 9], Subsection 9.4.2.1, Theorem 9.4.6. This distribution is absolutely continuous on the real line, but its density is explicitly known only for $\alpha=2$, when it becomes the Laplace distribution. The positive $N_{p}$-stable distribution has Laplace transform with $\beta>0$ and $\alpha\in(0,1]$:
@@
This is commonly known as the Mittag-Leffler distribution; see [23, Chapter 9], Subsection 9.4.2.2, Theorem 9.4.9. Also, in [13, Chapter 2], these were named Kovalenko distributions, see (3.2) or page 30. This distribution is also absolutely continuous, but its density is explicitly known only for $\alpha=1$ or $\alpha=0.5$. In the former case, this distribution is exponential, as in (13). In the latter case, for $\beta=1$, it has density
@@
These were the earliest and most studied examples. An entire chapter [23, Chapter 9] is devoted to these geometric stable distributions. See also a great survey [29] with large bibliography. Such geometric stable distributions were studied as far back as in 1984 in [26] and also in one early article by T. Kozubowski [28]. Further representation results for the Linnik and the Mittag-Leffler distributions are given in [27].
@@ 2.7. The infinitely divisible case
The following example is well known, see [17, Chapter I, Example 8.5; Chapter V, Example 13.2], [38, Section 8], [24, Section 6], Exampples 2 and 4, [7, Example 1]; [42, Example 3], and physical applications in [16, Section 8]. Here $X$ is a Gamma random variable with shape $1/k$ and certain rate parameter $\beta>0$. We remind the readers that it has Laplace transform
@@ 2.8. Commuting distributions
The key result is below. It was shown in other literature in particular cases, but not in this general form. For such results for or Gaussian $N$-stable random variables, see [13, Chapter 4], Theorem 4.6.1 and Corollary 4.6.1; [25, Theorem 1], [23, Chapter 8], Theorem 8.3.4 and references therein; [42, Theorem 2.1]. The property of commutative semigroups was also shown in [29, Remark 2.4]. This survey [29] is very comprehensive for 1990s. However, they claim that geometric family is unique among such explicit commuting families. We discuss in Section 2 that this is not true.
@@ 3.1. Introduction to the problem
The idea of the proof of Lemma 7 is to consider the behavior of the probability generating function of $N$ at $s=0$. From (2), which involves the behavior of the Laplace transform of $X$ at infinity. This approach is useful to disprove that $\varphi$ is a probability generating function in other cases. Such distribution $Q$ might have an atom at 0, that is, have $Q({0})=\mathbb{P}(X=0)>0$; and this happens if and only if $N$ has an atom at zero: $\mathbb{P}(N=0)>0$. By the way, the case of an atom at zero can be reduced to the case of no atom at zero. See [1, Chapter I], Section 12; [38, Section 9]; [16, Section 3], Definitions 3.1, 3.2; Theorem 3.2.
@@ 3.4. Parameterization
In the second case, we cannot have both expectation and convergence assumptions for $\nu_{p}$. The notation of $\nu_{p}$ is used in the following literature (the list is not exhaustive): [42, Section 2], (1) and (3), both expectation and convergence; [30, Section 1], only convergence; [24, page 304], they mentioned only expectation but clearly meant convergence as well, as seen from Theorem 2; [25, Section 2], only expectation, but later convergence is introduced; [30, Section 2] (3) and (4), only convergence; [13, Chapter 4], (6.32), only convergence. Sometimes, commutativity is required or derived in special cases.
The real reason behind the prevalence of this $p$-notation is that, historically, random stability study to a large extend (although not exclusively) was focused on geometric stability (that is, when $N=N_{p}$ is geometric from (11)). In this case, $pN_{p}$ converges to the standard exponential distribution as $p\downarrow 0$, and $\mathbb{E}[N_{p}]=1/p$. Approximations of sums of $N_{p}$ many independent identically distributed random variables as $p\downarrow 0$ (similar in spirit to the Central Limit Theorem) were extensively studied. A good exposition is in [23, Chapter 9].
@@ 4. Equivalent Definitions of Random Stability
There is another definition of random stable variables in the literature, one that is based on weak limits of scaled random sums. Just as stable and strictly stable distributions (in the classic sense) are often defined as weka limits of scaled sums of $N$ independent identically distributed random variables as $N\to\infty$, so can be done for random stable distributions. See [30, Section 2], and [38, Theorems 1, 2, 3].
@@
We see that $U_{j}$ belong to the strict domain of attraction: Recall our discussion above about defining strictly stable distributions as weak limits of scaled sums of independent identically distributed random variables. We stress this is different from classic $\alpha$-stable domains of attraction, where adding constants is allowed: Recall (6) and the discussion there. Strict domain of stability include scaling sums of independent indentically distributed random variables, but not adding constants. See also discussion in [30, Section 2].
@@ 5.3. Cyclic semigroup
$S(X)=\{1,a,a^{2},\ldots\}$ for some $a>1$. A construction was given in [7, Examples 2, 4] based on [9, Section 5] and [4] but in our view the proofs are not witten well, and they should be improved. Particular cases of $Q$ discussed in [4, Section 6] are as follows: The standard $P$-stable $Q$ for the following $P$: (a) $P(\{1\})=P(\{2\})=0.5$, with $\varphi(s)=(s+s^{2})/2$; (b) $P$ is the shifted geometric, with probability generating function $\varphi(s)=s^{2}/(4-3s)$, used in [2, Section 3] as a tool to construct Brownian motion on the Sierpinski gasket. The idea is to create an almost constant solution $\theta:(0,\infty)\to(0,\infty)$ to the functional equation $\theta(ax)\equiv b\theta(x)$, for fixed $a,b>0$. See [55] for additional information, also [46, Section 4].
@@ 5.5. Complete squares
$S(X)=\{1,4,9,16,\ldots\}$. This was discussed in [24, Section 4] for $L(u)=1/\mathrm{ch}(\sqrt{2u})$, where $\mathrm{ch}(x)=(e^{x}+e^{-x})/2$ is the hyperbolic cosine, with the inverse function $\mathrm{arcch}(s)=\ln(s+\sqrt{s^{2}-1})$.
@@
This equality in law follows from the self-similarity of the Brownian motion, see [20, Chapter 2]. We also refer the reader to [24, Section 3] with an interesting discussion about rational probability generating functions.
@@ 6.1. Construction
This branching process is described as follows, as outlined in [17, Chapter 5], Theorem 9.1, and [1, Chapter III], Sections 2 and 3, see also a short discussion in [51, Chapter V], Section 8. Start, as in discrete time, from $N(0)=1$. Fix a distribution $H$ on $\{0,2,3,\ldots\}$ called the generating distribution. This $H$ plays the role similar to the infinitesimal generator for diffusion processes. Each particle which exists at time $t$ is replaced during time interval $t,t+\triangle t]$ with small $\triangle t$, with probability $\triangle t$ with the number of particles distributed as $H$.
@@ 6.2. Yule and Neveu processes
For geometric distributions, $(N(t),\,t\geq 0)$ is called the Yule process, introduced by Yule in [54, Section 1] in 1925 for modeling evolution of the number of species as the pure birth process with birth rate $n$ at state $n$. See the discussion in [17, Chapter 5, Section 8] or [1, Chapter III], Section 5. For Sibuya distributions, this is called the Neveu process, introduced in an unpublished manuscript [44] by Neveu in relation to Generalized Random Energy Model (GREM) of Derrida. Links to Sherrington-Kirkpatrick model and GREM are discussed in [3, Chapter 6]. The exponential $\mathrm{Exp}(1)$ from (13) is $G_{p}$-stable for any $p$. However, there is no Sibyua-stable random variable, since the mean of the Sibyua distribution is zero. Properties of the Neveu process are discussed in [18]. They also considered generalizations with Sibuya distribution with an atom at zero in [18]. These branching processes have nonzero extinction probability.
@@ 6.3. Generating geometric distribution
Here we present a special case of geometric generating $H$. The full treatment is given in a recent article [40], with complicated formulas for mass functions and PGFs, using special functions. Our example is easily treatable and explicit. We also refer to another article [41] by the same authors, treating the case of Poisson $H$. This involves complicated special functions, too. Similar computations were done for [39, Theorem 1] by the same authors; but they consider the subcritical case $c<1$, with a.s. extinction; of course, this branching process does not have a limiting $X$. One possible application is to physics: the Bose-Einstein distribution have geometric as one possible limit, see [19, (9)], and this can be connected to the current process.
@@ 6.4. Theta-branching processes
In a recent article [34], they introduced a classification of 9 cases, see Section 4, with generalization of a Neveu process with Sibuya distribution. It has four parameters, including $\theta\in[-1,1]$. The case 4 is precisely the Sibuya distribution modified so that there is an atom at zero, and a nonzero extinction probability. For us, the case 3 is the most interesting, since it corresponds to the supercritical case, but with finite mean. The probability generating function $h$ of the generating distribution $H$ is, see [34, Section 7]:
@@
The probability generating function of the continuous-time branching process at time $t$ with $p=e^{-t}$, see [34, Section 4]: